OpenStudy (anonymous):

Let f(x)=x^3+x. If h is the inverse function of f, then h'(2) = The answer is 1/4 but I have no idea how. This is what i thought to do, and if someone could correct me it would be greatly appreciated. F(x)=x^3+x X=y^3+y 1=3y^2(y')+y' 1=y'(3y^2+1) 1/(3y^2+1)=y' where y=(x^3+x)

OpenStudy (anonymous):

sorry... i didn't mean that...

OpenStudy (anonymous):

ok... can you calculate f(1)...

OpenStudy (anonymous):

so f(1)=1^3+1 = 2

OpenStudy (anonymous):

ok since f(1)=2, that means \(\large f^{-1}(2)=1 \)

OpenStudy (anonymous):

now... what's f'(1)

OpenStudy (anonymous):

f'(1)=3(1)^2+1 = 4

OpenStudy (anonymous):

so since \(\large f'(1)=4 \) then \(\large f'(2)=\frac{1}{4} \)

OpenStudy (anonymous):

because the derivative is just the reciprocal...

OpenStudy (anonymous):